OpenStudy (anonymous):

Solve. (square root of 6)^8x = 216^x–3

OpenStudy (anonymous):

(sqrt 6)^(8x) = 216^(x - 3) sqrt 6 = 6^(1/2) and 216 = 6^3 (6^(1/2))^(8x) = (6^3)^(x - 3) Exponent rules. 6^(4x) = 6^(3x - 9) Take base 6 log of both sides. 4x = 3x - 9 x = -9